# Evolutionary Computation in the Era of Large Language Model: Survey and Roadmap

### v1

@@ II-B3 Algorithm Improvement
For an optimization model, if its constraints cannot be satisfied simultaneously, this model is called infeasible. Infeasible models are very common in practical applications, mainly because the model parameters are set incorrectly, or some conflicting new constraints are added. In order to diagnose the model and find the main constraints that cause the infeasibility, Chen et al. propose OptiChat [51], which uses LLMs to provide a natural language interface for non-expert users, making it more convenient to understand and repair infeasible optimization model problems. OptiChat can provide explanations of the optimization model, identify potential sources of infeasibility, and offer suggestions to make the model feasible through interactive conversations.
@@ III EA-enhanced LLM
This section delves into the realm of enhancing LLMs through the ingenious application of EA. The multifaceted exploration encompasses two primary dimensions: EA-based prompt engineering and EA-based LLM architecture search.
@@ III-A EA-based Prompt Engineering
Currently, evolutionary computation plays a role in two types of prompt engineering: discrete prompt optimization and continuous prompt optimization. The former considers meaningful textual instructions as prompts, while the latter uses vectors composed of continuous numerical values as prompts.
@@ III-A1 Textual Prompt Optimization
Guo et al. proposed an EA-based framework called EvoPrompt for discrete prompt optimization [20]. In EvoPrompt, the LLM emulates evolutionary operators using a fixed prompt and generates new candidate prompts through crossover and mutation based on the differences between parent prompts. The EA is used to guide the optimization process and retain the best prompts. EvoPrompt has been validated using two evolutionary methods: genetic algorithm and differential EA. In contrast to EvoPrompt’s initialization approach (hand-designed task-specific prompts), Fernando et al. introduced PromptBreeder, an automated approach that utilizes LLM to optimize prompts based on problem descriptions [21]. PromptBreeder employs an EA framework to automatically explore prompts in the problem domain and simultaneously evolves task prompts and mutation prompts, rather than using fixed prompts in EvoPrompt. Additionally, PromptBreeder adopts diverse mutation operators, including zeroth-order, first-order, and higher-order mutations, to effectively enhance the diversity of the prompt space. Another similar study to these two works is SPELL [22], which selects multiple parent prompts to generate ”meta-prompts” and provides them to LLM to generate descendant prompts. The selection of prompts is based on the fitness scores derived from the downstream task’s accuracy using a roulette wheel approach. Yang et al. argued that existing prompt engineering research primarily focuses on a single objective of model performance, whereas the quality of instructions should not be measured solely from a performance perspective; other detailed objectives such as instruction length and perplexity can be considered as well [52]. Similar to the aforementioned studies, Yang et al. proposed InstOptima, which utilizes the ChatGPT to simulate instruction mutation and crossover operations. Under multi-objective optimization, InstOptima employs the NSGA-II algorithm [53] for non-dominated sorting and obtains a set of excellent instructions (Pareto front) in terms of multiple objectives.
@@ III-A2 Gradient-Free Soft Prompt Optimization
In some open-source LLMs, users can only access the model through the model’s inference API and do not have access to the model parameters and gradient information. To optimize continuous prompts using limited samples in this scenario without relying on gradient-based optimization of model performance, Sun et al. proposed Black-Box Tuning (BBT) [54]. BBT represents the continuous prompts to be optimized as a vector to be optimized, and utilizes a fixed initialization vector. It then projects the prompt vector into a low-dimensional subspace using a random projection matrix to reduce the optimization difficulty. The Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES) algorithm [55] is used to sample new values for this vector in the subspace, and the loss value of the vector is queried from the LLM to modify the distribution for the next iteration of the EA’s search. This iterative process continues until the optimal solution for the subspace vector is obtained, which is then added to the fixed initialization vector and projected to obtain the final prompt vector. To further improve performance, Sun et al. proposed BBTv2 [56]. BBTv2 adds continuous prompt vectors not only at the input layer but also at each hidden layer in LLM, forming a deep prompt representation. This significantly increases the number of tunable parameters, which is advantageous for handling more challenging tasks. To optimize higher-dimensional deep prompts, BBTv2 employs a divide-and-conquer algorithm to alternately optimize prompt vectors at different levels, decomposing the original problem into multiple lower-dimensional subproblems. Additionally, the random projection matrix is no longer generated using a uniform distribution but is adapted to different models based on the statistical distribution of model word vectors or hidden states, better accommodating different models. Experimental results demonstrate that BBTv2 outperforms BBT on multiple comprehension tasks and models. To improve search efficiency and provide more fine-grained and multi-perspective evaluation feedback, Chai et al. replaced the method of reducing the space dimensionality by using a single random projection matrix in BBT with the Clip-Tuning approach [57]. Clip-Tuning applies Dropout sampling to the pre-trained model during inference, generating multiple sub-networks that can be viewed as projections of predictions for the samples in the original high-dimensional space. By blending rewards from multiple sub-network predictions, the search algorithm converges faster to the optimal solution. Similar to BBT, the EA in both of these studies is the CMA-ES.
@@
In addition to prompt engineering aimed at optimizing prompts, the prompt generation can have broader applications, such as data augmentation for training LLM or conducting jailbreak attacks on LLM to test security. The following will introduce how EA plays a role in these two tasks.
@@ III-A3 Prompt Generation for Data Augmentation
In contrast to the aforementioned approaches that focus on prompt engineering, Xu et al. employed prompt data augmentation using LLM and EA, referred to as Evol-Instruct, to enhance instruction data [58]. The framework uses a small number of manually written seed instructions as the initial dataset, treats LLM as an instruction evolver, iteratively evolves instructions based on different evolutionary prompts (e.g., adding constraints, deepening), and uses LLM to generate instruction responses and filterout ineffective instructions. The evolved instructions are then added to the training data, and the process is repeated iteratively to generate diverse and high-quality instruction sets. Experimental results demonstrate that Evol-Instruct improves the performance of instruction-based models on various downstream tasks, such as language modeling and text classification.
@@ III-A4 Prompt Generation for LLM Security
Another improvement of LLM is achieved by “jailbreak attacks” [59], which is a method to observe the security of LLM models. Specifically, alignment techniques ensure the safety and interpretability of model outputs by collecting human-annotated data and reinforcement learning methods [60], making the models generate responses that are more in line with human values and expectations. However, with properly designed problem prompts, aligned language models can still generate inappropriate responses, known as “jailbreak attacks”. Currently, handcrafted prompts for jailbreak attacks are difficult to automate and scale, while automatically generated prompts are often semantically incoherent and difficult to defend against semantics-based defenses [61]. Therefore, some studies use EA-based prompt engineering for automatic jailbreak attacks. Liu et al. models the jailbreak attack problem as an optimization problem and uses EAs to automatically optimize prompts [62]. It utilizes existing handcrafted jailbreak prompts to initialize the population and designs an adaptive function suitable for structured discrete data like text. In addition, AutoDAN adopts a hierarchical genetic algorithm to consider the optimization of prompts at the sentence and vocabulary levels, and designs a momentum word scoring mechanism to balance search ability and semantic coherence. Experimental results show that the jailbreak prompts generated by AutoDAN are semantically more coherent and effectively achieve jailbreak attacks. The research by Lapid et al. achieves a similar goal to AutoDAN [63].
@@ III-B EA-based LLM Architecture Search
With the increasing complexity and size of neural networks, manual design and optimization of architectures have become laborious and time-consuming tasks. EAs offer an effective approach to automate the search process and discover optimal architectures [64, 65]. By employing evolutionary operators such as mutation, crossover, and selection, these algorithms can generate a diverse set of candidate architectures. This exploration-exploitation trade-off allows for a comprehensive exploration of the architecture space while gradually converging towards promising solutions. Previously, EA-based Neural Architecture Search (NAS) methods were primarily applied to small-scale models and achieved promising results. Since the study by So et al. focusing on NAS for Transformers [66], EA-based NAS have started to be utilized on large-scale sequence models, including the currently popular diverse range of LLM architectures [^2] The papers reviewed in this subsection typically focus on designing NAS methods for pretrained language models (PLMs), without explicitly addressing the concept of LLMs. In our description, we do not differentiate between the concepts of LLMs and PLMs..
@@
Gao et al. used an NAS algorithm based on evolutionary search for the first time to discover a new universal LLM backbone from scratch, namely AutoBERT-Zero [23]. In AutoBERT-Zero, a well-designed search space is introduced, containing primitive math operations to explore novel attention structures in the intra-layer level, and leveraging convolution blocks as supplementary to attention in the inter-layer level. Additionally, the paper proposed the Operation-Priority evolution strategy which utilizes prior information of operations to flexibly balance exploration and exploitation during search. Furthermore, a training strategy called Bi-branch Weight-Sharing is designed to speed up model evaluation by initializing candidates with weights extracted from a super-net. Different from the more complex search space in AutoBERT-Zero [23], SuperShaper proposed by Ganesan et al. [24] focuses on searching the hidden unit dimensions of each Transformer layer. This is achieved by adding bottleneck matrices between each Transformer layer, thus enabling the variability of hidden dimensions across layers. By slicing the bottleneck matrices, different structured sub-networks are sampled from the super network, and the hidden dimensions between layers are determined by the slicing. The sub-network search process considers optimizing two objectives simultaneously, including perplexity and latency, and trained predictors are used to approximate them. Finally, an EA is used to search for the optimal hidden unit shape distribution across layers that simultaneously meets the requirements of accuracy and latency.
@@
Some methods try to achieve NAS on more complex search space, where key hyperparameters of each layer of Transformer are considered, such as hidden state dimension, number of attention heads, and feedforward network dimension. Yin et al. applied EA-based NAS methods for the first time to automatically optimize the architecture hyperparameters of LLMs [25]. The proposed AutoTinyBERT searches within the space of structures with identical layer depths and dimensions, thereby simplifying the search space from exponential to linear scale and greatly reducing the search complexity. In addition, the paper trained a big model SuperPLM with one-shot learning that contains all potential sub-structures. Therefore, when evaluating a specific structure, they do not need to train it from scratch, but can directly extract the corresponding sub-model from SuperPLM to serve as high-quality initializations for various latent structures. Yin et al. used EAs to search, and designed a sub-matrix extraction method to quickly extract different structural sub-models from SuperPLM for evaluating structure performance, selecting elite structures and mutating to generate new generations, repeating this process to search for the optimal structure. LiteTransformerSearch proposed by Javaheripi et al. [67] also allows the hyperparameters of each decoder layer, such as hidden size, number of heads and feedforward dimensions, to vary independently, creating a heterogeneous search space. The search space also includes other hyperparameters like number of layers and embedding dimensions. LiteTransformerSearch leverages the empirical observation that decoder parameter count has a high correlation with validation perplexity, establishing the first training-free low-cost proxy for Transformer architecture search - decoder parameter count. It uses EAs to sample candidate architectures from the search space based on this proxy, while also measuring hardware metrics directly on the target device. This enables LiteTransformerSearch to perform multi-objective NAS to obtain a Pareto frontier estimation that optimizes perplexity, latency and memory. And its effectiveness has been evaluated on two popular autoregressive Transformer backbones GPT-2 and Transformer-XL.
IV Integrated Synergy and Applications of LLM and EA

### v3

@@ II-C LLM Empowering EA with Other Capabilities
In addition to using LLM as an evolutionary operator or directly generating optimization algorithms, LLM can also provide assistance for EAs from other aspects. Chen et al. proposed OptiChat [53] to diagnose the infeasible model, identifying the main constraints that cause the infeasibility. Specifically, the infeasible model means that the constraints of an optimization algorithm cannot be satisfied simultaneously [54], which are very common in practical applications, mainly because the model parameters are set incorrectly, or some conflicting new constraints are added. OptiChat uses LLMs to provide a natural language interface for non-expert users, making it more convenient to understand and repair infeasible optimization model problems. The user provides the definition of the optimization model, including decision variables, parameters, constraints, and objective function, by using the Pyomo algebraic modeling language [55]. OptiChat can provide explanations of the optimization model, identify potential sources of infeasibility, and offer suggestions to make the model feasible through interactive conversations.
Besides, Maddigan et al. leveraged LLMs to provide explainability for results of EA, improving understanding for non-expert users [56]. They focused on using GP for nonlinear dimensionality reduction, where GP individuals are tree expressions that directly map each new low-dimensional feature to combinations of high-dimensional features. Specifically, a system is constructed for running GP-based nonlinear dimensionality reduction processes and visualizing results. LLMs are integrated into the system to develop a conversational interface to provide explanations, using prompts containing dataset and tree expression information to provide contextual knowledge. They also built a vector repository utilizing vector similarity search to provide related literature knowledge to the model and address knowledge gaps. Another recent study uses LLM to help decision makers understand and interpret the optimal solution set in multi-objective evolutionary optimization [57].
@@ III EA-enhanced LLM
This section delves into the emerging research of enhancing LLMs through the application of EA. The primary focus of this exploration is twofold: EA-based prompt engineering and EA-based LLM architecture search, as summarized in Tables III and IV.
@@ III-A EA-based Prompt Engineering
Black-box prompt engineering [58] enables adjusting prompts without requiring access to the underlying model’s parameters and gradients, making it particularly valuable for closed-source LLMs [59]. This advancement allows for effective optimization of closed-source models, overcoming previous limitations imposed by the unavailability of model parameters. Currently, evolutionary computation plays a role in two types of prompt engineering: discrete prompt optimization and continuous prompt optimization. The former considers meaningful textual instructions as prompts, as shown in Section III-A1, while the latter uses vectors composed of continuous numerical values as prompts, as shown in Section III-A2.
TABLE III: Summary of EA-based Prompt Engineering.
Type Method Used EA Ref.
GPS GA [60]
GrIPS Greedy search algorithm [61]
EvoPrompt GA, Differential EA [15]
Plum Metaheuristic algorithm [62]
Discrete PromptBreeder GA [63]
Prompt SPELL Any EA [64]
Optimization EoT prompting Any EA [65]
iPrompt Iteration similar to EA [66]
PhaseEvo Designed framework [67]
InstOptima NSGA-II [68]
EMO-Prompts NSGA-II, SMS-EMOA [69]
BBT CMA-ES [70]
BBTv2 CMA-ES [71]
Gradient-Free Soft Clip-Tuning CMA-ES [72]
Prompt Optimization Shen et al. CMA-ES [73]
BPT-VLM CMA-ES [74]
Fei et al. CMA-ES [75]
Prompt Generation for Evol-Instruct Mutation and selection [76]
Data Augmentation Sun et al. Any EA [77]
AutoDAN GA [78]
Prompt Generation Jailbreak Attacks GA [79]
for Security SMEA Any EA [80]
Shi et al. Any EA [81]
@@ III-A1 Textual Prompt Optimization
Before the emergence of ChatGPT, several research studies had already explored the use of metaheuristic algorithms, such as EAs, for prompt engineering in pretrained language models. Notable examples include Gradient-free Instructional Prompt Search (GrIPS) [61], which employed a greedy search strategy, and Genetic Prompt Search (GPS) [60], which utilized a Genetic Algorithm (GA) [82]. These studies typically employed EAs as the underlying search framework, with the LLM being responsible for generating and evaluating prompts. However, it is important to note that these investigations were limited in their scope and primarily focused on specific prompt engineering scenarios.
In order to fully unlock the potential of discrete optimization in the realm of black-box prompt optimization, Guo et al. proposed an EA-based discrete Prompt tuning framework (EvoPrompt) [15]. In EvoPrompt, the LLM emulates evolutionary operators using a fixed prompt and generates new candidate prompts through crossover and mutation based on the differences between parent prompts. The EA is used to guide the optimization process and retain the best prompts. EvoPrompt has been validated using two evolutionary methods: GA and Differential EA [83]. Pan et al. proposed a similar prompt optimization paradigm called Prompt Learning Using Metaheuristic (Plum) [62], which formulates prompt learning as a black-box non-convex discrete optimization problem, allowing various metaheuristic algorithms to be applied to help discover effective prompts. Beside the GA [82], this work implemented another five algorithms, including Hill Climbing, Simulated Annealing [84], Tabu Search [85], and Harmony Search [86]. Their experiments showed that Harmony Search was more efficient than GA, achieving better performance with fewer API calls.
In contrast to EvoPrompt’s initialization approach (hand-designed task-specific prompts), Fernando et al. introduced PromptBreeder, an automated approach that utilizes LLM to optimize prompts based on problem descriptions [63]. PromptBreeder employs an EA framework to automatically explore prompts in the problem domain and simultaneously evolves task prompts and mutation prompts, rather than using fixed prompts in EvoPrompt. Additionally, PromptBreeder adopts diverse mutation operators, including zeroth-order, first-order, and higher-order mutations, to more thoroughly explore the space of prompt strategies. Cui et al. proposed a prompt optimization framework called PhaseEvo that enables joint optimization of prompt instructions and examples [67]. The framework adopts a four-phase optimization strategy and designs five different evolutionary operators to achieve different goals in each phase, including global initialization phase (using Lamarckian operator or semantic mutation operator to generate initial population), local feedback mutation phase (using feedback mutation operator to accelerate convergence), global evolution mutation phase (using EDA operator and crossover operator to escape local optima), and local semantic mutation phase (using semantic mutation operator to find the global optimum). Additionally, PhaseEvo uses performance-based vectors and Hamming distance to evaluate candidate similarity instead of commonly used semantic similarity. Other similar approaches include Semantic Prompt Evolution based on a LLM (SPELL) [64], zero-shot EoT prompting [65], and interpretable autoprompting (iPrompt) [66].
Moreover, Yang et al. argued that the quality of instructions should not be measured solely from a performance perspective; other objectives such as instruction length and perplexity can be considered as well [68]. Similar to the aforementioned studies, Yang et al. proposed InstOptima, which utilizes the ChatGPT to simulate instruction mutation and crossover operations. Under multi-objective optimization, InstOptima employs the NSGA-II algorithm [87] for non-dominated sorting and obtains a set of excellent instructions (the Pareto front) in terms of multiple objectives. EMO-Prompts proposed by Baumann et al. also optimizes prompts from an evolutionary multi-objective optimization perspective [69]. It showcases an interesting scenario of finding prompts that cause the model to generate text containing two emotions.
@@ III-A2 Gradient-Free Soft Prompt Optimization
In some open-source LLMs, users can only access the model through the inference API and do not have access to the model parameters and gradient information. To optimize continuous prompts using limited samples in this scenario without relying on gradient-based optimization of model performance, Sun et al. proposed Black-Box Tuning (BBT) [70]. BBT represents the continuous prompts to be optimized as a vector with fixed initialization. It projects the prompt vector into a low-dimensional subspace using a random projection matrix to reduce the optimization difficulty. The CMA-ES is used to sample new values for this vector in the subspace, and the loss value of the vector is queried from the LLM to modify the distribution for the next iteration of CMA-ES. This iterative process continues until the stop condition is met, where the obtained subspace vector is then added to the fixed initialization vector and projected to obtain the final prompt vector.
@@
Several studies aimed to improving the BBT. Sun et al. proposed BBTv2 [71]. BBTv2 adds continuous prompt vectors not only at the input layer but also at each hidden layer in LLM, forming a deep prompt representation. This significantly increases the number of tunable parameters, which is advantageous for handling more challenging tasks. To optimize higher-dimensional deep prompts, BBTv2 employs a divide-and-conquer algorithm to alternately optimize prompt vectors at different levels, decomposing the original problem into multiple lower-dimensional subproblems. Additionally, the random projection matrix is no longer generated using a uniform distribution but is adapted to different models based on the statistical distribution of model word vectors or hidden states, better accommodating different models. To improve search efficiency and provide more fine-grained and multi-perspective evaluation feedback, Chai et al. replaced the method of reducing the space dimensionality by using a single random projection matrix in BBT with the Clip-Tuning approach [72]. Clip-Tuning applies Dropout sampling to the pre-trained model during inference, generating multiple sub-networks that can be viewed as projections of predictions for the samples in the original high-dimensional space. By blending rewards from multiple sub-network predictions, the search algorithm converges faster to the optimal solution. Shen et al. learned the complete distribution over soft prompts, rather than just considering a point estimate like BBT [73], in order to achieve uncertainty quantification of the predictive results. They used variational inference and ensemble learning to learn and approximate the posterior distribution over soft prompts. Similar to BBT, the EA in these studies is the CMA-ES. Some other studies focused on the vision-language models [74, 75], which are not detailed here.
@@
In addition to prompt engineering aimed at optimizing prompts, the prompt generation can have broader applications, such as data augmentation for training LLM or conducting jailbreak attacks on LLM to test security. Sections III-A3 and III-A4 will introduce how EA plays a role in these two tasks.
@@ III-A3 Prompt Generation for Data Augmentation
In contrast to the aforementioned approaches that purely focus on prompt optimization, Xu et al. employed prompt data augmentation using LLM and EA, referred to as Evol-Instruct, to enhance instruction data [76]. The framework uses a small number of manually written seed instructions as the initial dataset, treats LLM as an instruction evolver, iteratively evolves instructions based on different evolutionary prompts (e.g., adding constraints, deepening), and uses LLM to generate instruction responses and filter out ineffective instructions. The evolved instructions are then added to the training data, and the process is repeated iteratively to generate diverse and high-quality instruction sets. Experimental results demonstrate that Evol-Instruct improves the performance of instruction-based models on various downstream tasks, such as language modeling and text classification. Sun et al. adopted a similar approach to build high-quality domain-specific instruction data, and designed a multidimensional quality evaluation method to assess the generated data [77].
@@ III-A4 Prompt Generation for LLM Security
Another improvement of LLM focuses on model security. Specifically, alignment techniques ensure the safety and interpretability of model outputs by collecting human-annotated data and reinforcement learning methods [88], making the models generate responses that are more in line with human values and expectations. However, with properly designed problem prompts, known as “jailbreak attacks”, aligned LLMs can still generate inappropriate responses. Currently, handcrafted prompts for jailbreak attacks are difficult to automate and scale, while automatically generated prompts are often semantically incoherent and difficult to defend against semantics-based defenses [89]. Therefore, some studies use EA-based prompt engineering for automatic jailbreak attacks. Liu et al. model the jailbreak attack problem as an optimization problem and use EAs to automatically optimize prompts [78], namely Automatically generating DAN-series-like jailbreak prompts (AutoDAN). It utilizes existing handcrafted jailbreak prompts to initialize the population and designs an adaptive function suitable for structured discrete data like text. In addition, AutoDAN adopts a hierarchical GA to consider the optimization of prompts at the sentence and vocabulary levels, and designs a momentum word scoring mechanism to balance search ability and semantic coherence. The research by Lapid et al. achieves a similar goal to AutoDAN [79]. Zou et al. noted the importance of system messages in LLMs jailbreaking, as well as the transferability of jailbreaking prompts under different system messages. They proposed the System Message Evolutionary Algorithm (SMEA) [80] to search for optimized system messages with stronger resistance against jailbreaking attacks. Another study focused on testing the robustness of reliability detection systems for LLM-generated texts [81]. They utilized the protected auxiliary model ChatGPT to generate word substitution candidates, and performed query-based word substitution attacks based on EA. The search involved instructional prompts to change the generation style and make the detection system difficult to detect, which is essentially a prompt optimization process for adversarial purposes.
Based on the discussions above, EA-based prompt engineering has played a significant role in LLM. Researchers have used EAs as a search framework, combined with prompt generation and evaluation in both discrete and continuous prompt optimization. Additionally, EAs have also been applied to data augmentation and LLM security, further expanding the scope of prompt generation applications. However, current methods still face challenges such as the selection of initial populations, expanding search spaces, and method stability. Addressing these challenges will require better strategies for selecting initial populations, efficient search algorithms, and a deeper understanding of LLMs. Therefore, EA-based prompt optimization provides powerful tools and methods for improving LLM performance and expanding its applications. With further research and advancements, we can expect breakthroughs and innovations that will contribute to the progress of prompt optimization in LLM.
@@ III-B EA-based LLM Architecture Search
TABLE IV: Summary of EA-based LLM Architecture Search Methods
Method Main Task Used EA Evaluated LLM Ref.
AutoBERT-Zero Discover new universal LLM backbone from scratch Any EA BERT [16]
SuperShaper Search hidden unit dimensions in each Transformer layer Any EA BERT [90]
AutoTinyBERT Automatically optimize LLM architecture hyperparameters Any EA BERT [91]
LiteTransformerSearch Independently vary hyperparameters of each decoder layer Any EA GPT-2 [92]
Klein et al. Various LLM architecture optimization methods Multi-objective EA BERT [93]
Choong et al. Obtain specialized models optimized for specific tasks MO-MFEA M2M100-418M, ResNet-18 [94]
@@
Prompt engineering goes beyond the realm of LLM models by optimizing the input format to enhance the quality of model outputs. Differing from prompt engineering, another approach known as LLM architecture search focuses on directly optimizing the architecture of LLM models to achieve superior performance and lighter LLM models. With the increasing complexity and size of neural networks, manual design and optimization of architectures have become laborious and time-consuming tasks. EAs offer an effective approach to automate the search process and discover promising architectures [95, 96]. By employing evolutionary operators such as mutation, crossover, and selection, these algorithms can generate a diverse set of candidate architectures. This exploration-exploitation trade-off allows for a comprehensive exploration of the architecture space while gradually converging towards promising solutions. Previously, EA-based NAS methods were primarily applied to small-scale models and achieved promising results. Since the study by So et al. focusing on NAS for Transformers [97], EA-based NAS has been utilized on diverse large-scale models [^3] The papers reviewed in this subsection mainly concentrate on designing NAS methods for pretrained language models (PLMs) without explicitly distinguishing between the concepts of LLMs and PLMs. as shown in Table IV.
@@
The first work, using an NAS algorithm based on evolutionary search is AutoBERT-Zero [16], to discover a new universal LLM backbone from scratch. In AutoBERT-Zero, a well-designed search space is introduced, containing primitive math operations to explore novel attention structures in the intra-layer level, and leveraging convolution blocks as supplementary to attention in the inter-layer level. Additionally, this work proposed the Operation-Priority evolution strategy which utilizes prior information of operations to flexibly balance exploration and exploitation during search. Furthermore, a training strategy called Bi-branch Weight-Sharing is designed to speed up model evaluation by initializing candidates with weights extracted from a super-net. Different from the more complex search space in AutoBERT-Zero [16], SuperShaper proposed by Ganesan et al. [90] focuses on searching the hidden unit dimensions of each Transformer layer. This is achieved by adding bottleneck matrices between each Transformer layer, thus enabling the variability of hidden dimensions across layers. By slicing the bottleneck matrices, different structured sub-networks are sampled from the super network, and the hidden dimensions between layers are determined by the slicing. The sub-network search process considers optimizing two objectives simultaneously, including perplexity and latency, and trained predictors are used to approximate them. Finally, an EA is used to search for the optimal hidden unit shape distribution across layers that simultaneously meets the requirements of accuracy and latency.
@@
Some methods try to achieve NAS on more complex search space, where key hyperparameters of each layer of Transformer are considered, such as hidden state dimension, number of attention heads, and feedforward network dimension. Yin et al. applied EA-based NAS methods for the first time to automatically optimize the architecture hyperparameters of LLMs [91]. The proposed AutoTinyBERT searches within the space of structures with identical layer depths and dimensions, thereby simplifying the search space from exponential to linear scale and greatly reducing the search complexity. In addition, this work trained a big model SuperPLM with one-shot learning that contains all potential sub-structures. Therefore, when evaluating a specific structure, they do not need to train it from scratch, but can directly extract the corresponding sub-model from SuperPLM to serve as high-quality initializations for various latent structures. Yin et al. used EAs to search, and designed a sub-matrix extraction method to quickly extract different structural sub-models from SuperPLM for evaluating structure performance, selecting elite structures and mutating to generate new generations, repeating this process to search for the optimal structure. LiteTransformerSearch proposed by Javaheripi et al. [92] also allows the hyperparameters of each decoder layer, such as hidden size, number of heads and feedforward dimensions, to vary independently, creating a heterogeneous search space. The search space also includes other hyperparameters like number of layers and embedding dimensions. LiteTransformerSearch leverages the empirical observation that decoder parameter count has a high correlation with validation perplexity, establishing the first training-free low-cost proxy for Transformer architecture search - decoder parameter count. It uses EAs to sample candidate architectures from the search space based on this proxy, while also measuring hardware metrics directly on the target device. This enables LiteTransformerSearch to perform multi-objective NAS to obtain a Pareto frontier estimation that optimizes perplexity, latency and memory. And its effectiveness has been evaluated on two popular autoregressive Transformer backbones GPT-2 and Transformer-XL. Similarly, Klein et al. also conduct NAS in multi-objective manner [93], which discover multiple subnetworks of LLMs that balance model performance and size.
Different from the above methods, Choong et al. did not directly study the NAS task, but proposed the concept of “Set of Sets” [94], which refers to a set of models that can simultaneously meet multiple task settings and resource constraints. They studied how to obtain a set of smaller-scale models optimized for specific tasks from LLMs through multi-objective multi-task EAs (such as the MO-MFEA algorithm [98, 99]). Among them, the LLM plays the role of a general-purpose basic model. Experimental results show that the specialized models obtained in this way can achieve better performance or greater compression rates than the original large model in various application fields and neural network architectures.
In conclusion, EAs have shown great potential in assisting LLM architecture search. By leveraging EAs, researchers can automate the process of optimizing LLM architectures, which is otherwise laborious and time-consuming. EAs enable the generation of diverse candidate architectures through evolutionary operators such as mutation, crossover, and selection, striking a balance between exploration and exploitation. Several studies have successfully applied EA-based NAS methods to different aspects of LLM architecture, including discovering new universal LLM backbones, optimizing hidden unit dimensions, and tuning hyperparameters. However, there are still some challenges remaining, such as the high time consumption and the limited generalization ability. Future research should focus on addressing these challenges and exploring more effective LLM architecture search approaches.
III-C EA Empowering LLM for Other Enhanced Capabilities
In addition to enhancing the performance of LLM through NAS, some researchers have utilized the search capability of EA to assist in improving other aspects of LLM.
Before the emergence of LLM, Kim et al. have studied the Length-Adaptive Transformer model [100], which can automatically adjust the sequence length according to different computational resource constraints during inference. They proposed LengthDrop technique, which allows the model to randomly reduce the sequence length of each layer during training, making the trained model more robust to changes in sequence length during inference. On the trained model, an EA is used to continuously optimize the population of length configurations to maximize accuracy under various computational budgets. During inference, the corresponding optimal length configuration is directly used without retraining the model to achieve efficient inference. Jiang et al. studied how to deploy generative inference services for large foundation models in a heterogeneous distributed environment to reduce the costs of centralized data centers [101]. Their research allows each pipeline parallel stage to be assigned with different numbers of Transformer layers, while also allowing each stage to set different degrees of tensor model parallelism. They used a GA with operations of merging, splitting, and swapping different pipeline groups. By combining with a dynamic programming algorithm to evaluate the costs of each scheme, they searched for the optimal solutions of pipeline group allocation, GPU device allocation within each pipeline group, and layer allocation for each pipeline stage. Ding et al. studied how to extend the context window of LLMs to 2048k tokens [102]. They discovered two forms of non-uniformities in the positional encoding - across different dimensions and token positions. Through an EA, they searched for the optimal rescaling factors for each dimension and initial positions. This allowed different dimensions and positions to be rescaled with different degrees of interpolation or extrapolation. It achieved better preservation of the original positional information compared to existing methods that uniformly handled all dimensions. Akiba et al. proposed a model merging approach that utilizes CMA-ES algorithm to optimize merged performance in both parameter space and data flow space [103], which can automatically discover effective combinations of diverse models from different domains. Li et al. propose to leverage a small model to store domain knowledge in order to assist LLMs in tackling the lack of domain-specific knowledge [104]. To learn and apply the knowledge, the small model undergoes domain-specific pretraining, knowledge instruction tuning, and Bayesian optimization, where CMA-ES is used to find soft prompts that optimizes the consistency between the outputs of the two models. This allows the small model to better satisfy the needs of the large model in downstream tasks. Moreover, the studies of self-evolution in LLMs have also widely adopted the ideas of EAs, which has been discussed in [105].
IV Applications Driven by
Integrated Synergy of LLM and EA